Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-04-23

Observers: Sandra Benitez, Joe Polshaw
Support: Stefan Taubenberger
TAT on duty: Morgan Fraser, Nancy Elias-Rosa, Joe Anderson

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 2 m/s; temp. 16.2 C; humid. 21%; Clear. 
middle: wind 0.5 m/s; temp. 14.8 C: humid. 12%; Clear.
end: wind 0.6 m/s; temp. 14.6; humid. 11%; Clear

Clear night, with seeing averaging around 1" or slightly less. There were some fluctuations +/- 0.4" but nothing major. Photometric.

--------------------------------------------------------------

INSTRUMENTS: EFOSC2, SOFI (Flats were taken on 2014-04-21 for SOFI)

CALIBRATIONS

Afternoon:
-- Bias
-- EFOSC flats
-- EFOSC arcs

Evening:
-- Sky flats: UB

Morning:
-- EFOSC arcs

--------------------------------------------------------------

OBSERVATIONS

Target                   Type                         UT-Start         Filter/Grism           Slit (")         Seeing (")      Notes

*** EFOSC2 ***

SN2013ai              Follow. (IMAG)           23:29              VRi                        N/A             0.75              (1)

*** Change to SOFI ***

LSQ14an              Follow. (IMAG)            00:53                JHKs                      N/A             0.75             (2)
CSS140421...       Follow. (IMAG)            02:12                JHKs                      N/A              1                 (3)              
LSQ14asn             Follow. (IMAG)            02:52                JHKs                      N/A             0.6              (4)

*** Change to EFOSC2 ***

CSS140421...       Follow. (SPEC)            05:10               Gr11/16                  1                 0.6               (5)
CSS140421...       Follow. (IMAG)           06:43               UBVRi                      N/A              0.8              (6)
LSQ14asn            Follow. (SPEC)            07:13                Gr11                        1                 1.0              (7)
EG274                 STD. (SPEC)                08:40                 Gr11/13/16            1                1.0              (8)
LSQ14bef            Class.                         08:58                Gr13                         1              1.0               (9)                      
PG1657               STD (PHOT)                09:25                UBVRi                     N/A             1.0                -
SN2009ip            Follow. (PHOT)            09:35               VRi                        N/A              1.0               (10)
Marka                  STD (PHOT)                10:00               UBVRi                     N/A             1.0                -


(1) 6*150s in each band. The DIMM eeing increased to ~1.2 for a very short while in the middle of the OB, and decreased down to 0.6" by the end.
(2) Some delay to the start due to instrument change and image analysis. 
(3) CSS140421-142042+031602. We did some extra imaging after the initial OB as we waited for LSQ14asn (the next target) to decrease in elevation (it was right at the zenith at this point). We had planned to observe a standard star for the illumination mask in this period but were informed not to earlier in the night due to problems with the OB. There was also some fluctuation in the seeing during the first OB, between 1" and 1.5". There is a clear detection in each band.
(4) The stars in the first few J-band images may be *slightly* elongated due to the high elevation. After a quick look they appear to be fine (most definitely useable!). All H and Ks images should be fine.
(5) CSS140421-142042+031602. We lost some time here due to the instrument change and image analysis. 40mins gr11 + gr16. Forgot to fill in the "Name" field in the OB, so it may appear on the archive as "No Name". The next OB is imaging of same object so should be clear which files to download from the archive. Some fluctuations in the DIMM seeing, from 0.6" to 1.4".
(6) CSS140421-142042+031602. Looks bright in the images (esp. UB bands)
(7) The exposure time has been reduced to 2*2400s since there was a chance the target would be too low by the end of the observation. High airmass towards the end, with some fluctuation in DIMM seeing (between 0.5" and 1.5").
(8) Standard star using both 1" and 10" slit to check throughput. ** We loaded the OB for the wide slit, but it failed since we did not select slit 10" in the instrument set-up. So we skipped over the slit 10" sections of the OB and only did slit 1". We will include slit 10" in the instrument set-up for the next sub-run **
(9) Not sure what this one is...
(10) Ran out of targets - SN2009ip was the only target we could observe (except for more standards!) so we did some quick photometry (2*200s Vi, 3*100s R).


Comments